Calculate the standard entropy change for the following reaction at 25C you will have to look up the values for entropy change of the reactans and products

2Al(s)+3ZnO(s)->Al2O3(s)+3Zn(s)
Answer #1

Given reaction is

2Al(s)+3ZnO(s)-->Al2O3(s)+3Zn(s)

ΔS rxn = sum of ΔS (products) - sum of ΔS (reactants)

the standard values of entropy change are

ΔS Al2O3(s) = 50.99 J/K mol

ΔS ZnO = 43.9 J/K mol

ΔS Al = 28.3 J/K mol

ΔS Zn = 41.6 J/K mol

ΔSrxn = [ΔSAl2O3 + 3 X ΔSZn] -[2X ΔSAl + 3 X ΔSZnO]

Putting values

ΔSrxn = [50.99 + 3 X 41.6] -[2 X 28.3 + 3 X 43.9] = 175.79 - 188.3 = -12.51J /K mole